In January 2021 Vega Corporation purchased a patent at a cost of $215,000. Legal and filing fees of $52,000 were paid to acquire the patent. The company estimated a 10-year useful life for the patent and uses the straight-line amortization method for all intangible assets. In January 2024, Vega spent $29,000 in legal fees for an unsuccessful defense of the patent and the patent is no longer usable. The amount charged to income (expense and loss) in 2024 related to the patent should be:

Multiple Choice

  • $215,900.

  • $215,000.

  • $ 55,700.

  • $ 29,000.
Answer #1

Cost of patent = 215,000+52,000 = 267,000

Book value of patent on Jan 2024

= 267,000 - (267,000/10 * 3 years)

= 267,000 - 80,100

= 186,900

Amount charged to income = 186,900 + 29,000

= 215,900

Option A